40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3048     Entry time: Thu Jun 3 22:33:31 2010
Author: valera 
Type: Summary 
Category: CDS 
Subject: simulated plant work 

 I put matlab files and a summary into the 40m wiki for the fitting of the 40m Optickle transfer functions and generating digital filters for the simulated plant:

http://lhocds.ligo-wa.caltech.edu:8000/40m/Generating_DOF-%3EPD_digital_filters_based_on_Optickle_modeling

The filters are not loaded yet. Joe and Alex will make a rcg code to make a matrix of filters (currently 5x15=75 elements) which will enable the simulated plant tf's.

Joe and I tried to put a signal through the DARM loop but the signal was not going through the memory location in the scx part of the simulated plant.

 

Edit by Joe:

I was able to track it down to the spx model not running properly.  It needed the Burt Restore flag set to 1.  I hadn't done that since the last rebuild, so it wasn't actually calculating anything until I flipped that flag.  The data is now circulating all the way around.  If I turn on the final input (the same one with the initial 1.0 offset), the data circulates completely around and starts integrating up.  So the loop has been closed, just without all the correct filters in.

ELOG V3.1.3-